# We Have a Package for You! A Comprehensive Analysis of Package Hallucinations by Code Generating LLMs

### v2

@@ 4.2 Code Generation
Model Selection. For our experiments, we chose the top-ranked models from the EvalPlus leaderboard (as of January 2024) [45]. While creating our list, we disregarded the fine-tuned versions that were ranked below their corresponding foundational models, and only selected one fine-tuned version of the same foundational model of the same parameter size [45]. EvalPlus maintains a ranking of the top performing LLMs for code correctness according to a rigorous code synthesis evaluation framework. Our goal was to include a mix of top performing base models and a few of the best performing fine-tuned variants. We also included the GPT series of models (GPT-3.5, GPT-4, and GPT-4 Turbo) in our experiments, which currently hold the top rankings on the leaderboard. GPT models are widely considered as SOTA in terms of code generation models at the time of writing and add value to our experiments as representative commercial models. The models were not modified or altered in any way prior to testing; they are strictly “off-the-shelf.” Table 1 provides a complete list of the models that we tested in our experiments.
@@
Language Selection. In our experiments, we focus on two of the most popular programming languages, JavaScript and Python. These languages were chosen for their overall popularity (#1 and #2 according to GitHub’s 2023 Octoverse report [15]) and their reliance on open-source repositories for package management. Other popular programming languages like Java, C, or C++ do not rely on a centralized, open-source repository as Python and JavaScript do, which is a key component of this vulnerability. The open-source package repositories for these languages, npm and PyPI, represent ecosystems of 5.1 million and 573 thousand packages, respectively [72]. Of the 16 total models tested (see Table 1), fourteen were tested for both Python and JavaScript, while two fine-tuned Python-specific models, WizardCoder-Python and CodeLlama-Python, were only tested for Python.
Testing Environment. To create a fair and uniform testing environment, we utilize the text-generation-webui tool for LLMs, which supports most open-source LLMs and offers both a GUI and API for local and offline use [59]. It allows precise parameter control, ensuring consistent model evaluation during testing. We employed the GPTQ quantization method on these models, which reduces parameter precision to boost inference speed and lower memory use without significantly impacting performance. The GPTQ quantization method, specifically, uses a one-shot weight quantization method based on approximate second-order information that has a negligible effect on the accuracy of models, making it an ideal choice [13]. Additionally, quantized models better simulate the performance that a typical user can expect when running models on commercial grade hardware, making them more accessible and practical for everyday use.
For testing uniformity, we use the same parameters and quantization precision for all open-source models, which are summarized in Appendix C, along with the computing environment used. To generate code for our analysis, we query each of the LLMs (Table 1) with prompts from the two datasets using the Python requests packages to interact with the text-generation-webui API. The message to the API contains a system message with specific instructions for the model and the prompt itself. An overview of the process, including system messages used during each step, are detailed in Appendix B. This results in a total of 19,200 code samples per model (16 Python tests + 14 JavaScript tests * 19,200 = 576,000 total code samples), which are further analyzed to determine which packages are required to execute the generated code.
@@ 4.3 Detection Methodology and Heuristics
Detection of hallucinated packages from LLM outputs or code samples is non-trivial. In short, simply parsing the code for “import” or “require” is not useful as the arguments in those statements refer to modules and not packages. There is no way to definitively determine required packages from a code snippet alone. A detailed explanation of this problem can be found in Appendix G. To solve this problem, we employ the following three heuristics to determine/identify package names in the generated code:
@@
Heuristic 1. As part of our first heuristic, we parse the generated Python and JavaScript code for “pip install” and “npm install” commands, respectively. These commands look for the specified package in the PyPI/npm repository, resolve its dependencies, and install everything in the current Python/JavaScript environment to ensure that future module requests will work. This is the most straightforward heuristic for detecting package names (and thus, hallucinations), as it involves explicit commands from the code generation model for package download/installation. This is significant because if the referenced hallucinated package was indeed used by an adversary to execute a package confusion attack, it could immediately trigger download/install of the malicious code in the package. Note that we did not directly ask the model to provide these commands, but allowed them to occur naturally during the generation process. As such, we observed that these instances (“pip install” and “npm install”) occur for $7\%$ of the total outputs.
@@
Heuristic 2. For the second heuristic, each generated code sample is used as an input into the same model that generated it. The model is then prompted for a list of packages that would be required to run the given code. Our intuition is to mimic an actual user/developer who is using LLMs for code generation. If the user gets an error due to an uninstalled package when attempting to execute the generated code, they could query the model for the correct package to install. We wanted to replicate this intuitive process to identify package names required by the generated code.
@@
Heuristic 3. As the third heuristic, we re-use the original prompt used to generate the code sample as an input to the model and ask the model to output package names that would be required to accomplish this coding task. Similar to the previous heuristic, this process of extracting package names simulates another approach users would take to obtain package names from the model that generated the code, if the required packages were not mentioned in explicit “pip install” and “npm install” commands.
@@
Once each model provides specific package names (through the three heuristics outlined above), we simply compare each package name to a master list of package names acquired from PyPI and npm, respectively (each list is as of 10 January, 2024). If a package name is not on the master list it is considered a hallucination. We acknowledge the possibility that the master list of packages obtained from the package repositories has already been contaminated with malicious hallucinated packages. It is not possible to guarantee that the master list actually represents the ground truth of valid packages, however, the presence of hallucinated packages already in the master list would actually produce fewer hallucinations, therefore our results represent a lower bound of hallucination rate.
@@ 5 Evaluation Results
In this section, we present the results of our experimental analysis related to RQ1 - RQ4. After using both Python and JavaScript for RQ1, for RQs 2 through 4, we focus our analysis only on the Python programming language, a subset of the original models tested, and randomly sampled subsets of our original datasets. Given the consistent results that we were able to obtain across both languages in RQ1, this narrowed scope of discussion does not compromise the generalizability of the conclusions and allows for a deeper analysis of package hallucinations in a controlled setting. We selected GPT-4 Turbo, GPT-3.5, CodeLlama 7B, and DeepSeek 6.7B for in-depth analysis, representing the best-performing and most popular open-source models.
@@ 5.1 Prevalence of Package Hallucinations (RQ1)
In our first experiment, our goal was to quantify the prevalence of package hallucinations across different models by generating and analyzing a large number of code samples. We conducted 30 tests (using 16 models for Python and 14 models for JavaScript, as outlined in Table 1) producing a combined 576,000 code samples using both the Stack Overflow and LLM-generated datasets (Section 4.1). Each code sample was evaluated for hallucinations according to the heuristics defined in Section 4.3, which include parsing the generated code and prompting the model for packages twice per code sample, for a total of 1,152,000 package prompts across all tests. To measure LLMs’ propensity to produce hallucinated packages during code generation we use the package hallucination rate metric, which can be expressed as a simple ratio of the number of hallucinated packages to the total number of recommended packages. The full set of hallucination results for all models, covering both Python and JavaScript, are presented in Appendix E.
@@
These 30 tests generated a total of 2.23 million packages in response to our prompts, of which 440,445 (19.7%) were determined to be hallucinations, including 205,474 unique non-existent packages (i.e. packages that do not exist in PyPI or npm repositories and were distinct entries in the hallucination count, irrespective of their multiple occurrences). Our results for GPT-3.5 (5.76%) and GPT-4 (4.05%) differ significantly from previous work on package hallucinations [35], which found hallucination rates 4-6 times greater (24.2% and 22.2%, respectively) for those specific models. GPT-series models were found 4 times less likely to generate hallucinated packages compared to open-source models, with a 5.2% hallucination rate compared to 21.7%. GPT-4 Turbo resulted in the lowest overall hallucination rate at 3.59%, while DeepSeek 1B had the best hallucination rate among the open-source models at 13.63%. Python code resulted in fewer hallucinations than JavaScript (15.8% on average compared to 21.3% for JavaScript). Despite the difference in the hallucination rate between the two languages, there is a linear relationship between the results (as shown in Figure 14 in the appendix), which demonstrates that the propensity of a model to hallucinate is positively correlated between programming languages. The total hallucination rates for each evaluated model are presented in Figure 2. For a comprehensive breakdown of the results for each language, please refer to Appendix E. These results provide strong evidence that package hallucinations are a pervasive issue across all code generating LLMs.
RQ1 Summary: Package hallucinations were found to be pervasive phenomenon across all 16 models tested. Commercial models hallucinated 4x less compared to open-source models. Python code resulted in a lower hallucination rate compared to JavaScript.
@@ 5.2 Impact of Model Settings (RQ2)
Effect of Temperature Settings. The temperature setting in LLMs adjusts the randomness of the generated responses, where a lower temperature results in more predictable and deterministic outputs, and a higher temperature increases creativity and diversity in the responses (Section F.2). We varied this setting for each model between the minimum and maximum allowed values and observed the change in hallucination rate (the maximum temperature for GPT series models is limited to 2, while the open-source models can be set up to 5). All models exhibited a clear increase in hallucination rate as temperature value increases, with the effect becoming severe at maximum values. The OpenAI models, as shown in Figure 3, displayed only a slight increase in hallucination rate between temperatures 0 and 1, which then increased sharply between 1 and 2. Most notably, GPT-4 resulted in a hallucination rate (8.9%) nearly 4x lower than GPT 3.5 (31.8%) at its maximum temperature. At the highest temperature values, open-source models start to generate more hallucinated packages than valid packages. Most LLMs operate at a default temperature in the range of 0.7 to 1, however, our results indicate that a lower temperature value can reduce package hallucinations, with the optimal value varying per model. A lower temperature also yields more deterministic responses, presenting a trade-off between risk of hallucination and creativity. Therefore, selecting the appropriate temperature value is not a straightforward decision.
@@
Effect of Decoding Strategies. Next, we adjusted several decoding parameters (top-$p$, top-$k$, and min-$p$ values) to reduce the chances of a low probability token being selected as a potential package, with the intuition that lower probability tokens correspond to higher probabilities of hallucination in this context. Below is a summary of the parameters and the values we modified. We evaluated each listed value in isolation, followed by a combined evaluation of the values highlighted in bold, resulting in a total of 10 tests. Note that top-$k$ and min-$p$ were only tested for DeepSeek and CodeLlama, as those values are not modifiable through the OpenAI API.
@@
Varying the decoding values induced a slight increase (1.16% on average) in hallucination rate for all four models across all values tested. As we will expand on in RQ3, package hallucinations are often persistently repeated across many iterations.
@@
Recency of Subject Matter. As described in Section 4.1, we separated our coding prompts into two temporal datasets to evaluate if hallucination rate was correlated with topics/packages that emerged after the model was trained. A lower difference between the rates of recent and all-time prompts indicates better performance in handling questions that fall outside the model’s pre-training data and therefore a more generalizable model. The models we tested were shown to be more likely to generate a package hallucination when responding to prompts that deal with more recent topics. This difference resulted in a 10% higher hallucination rate on average for older data vs. more recent data.
Overall, all 16 Python models we evaluated demonstrated a higher hallucination rate when being prompted about questions or packages that were popular within the past year (Figure 4). These higher rates are at least partly due to inherent limitations and training costs of modern LLMs. As noted in the OpenAI GPT-4 technical report [1], LLMs cannot update themselves with new information post-release and have no knowledge of the world beyond their training data cut-off date. Although fine-tuning can enhance specific tasks, it does not generally improve the model’s overall knowledge of the world. The massive cost of training LLMs from scratch, evidenced by the 1,400,000 GPU hours (220 years) required to train the 12 CodeLlama models, makes continuously updating pre-training data impractically expensive [66]. This cost, along with steadily increasing model sizes and training times, poses a significant barrier to reducing package hallucinations for advanced coding prompts and packages.
@@ 5.3 Model Behaviors (RQ3)
Frequency of Repeated Hallucinations. To determine whether hallucinations are random error or repeatable phenomena, this test focuses on the persistence of hallucinations within a model. We randomly sampled 500 prompts that generated package hallucinations during our initial testing, then repeated those queries 10 times per prompt. Of those 10 queries we recorded how many times the original hallucinated package was regenerated. Our analysis reveals an unexpected dichotomy when repeatedly querying a model with the same prompt that generated a hallucination: 43% of hallucinated packages were repeated in all 10 queries, while 39% did not repeat at all across the 10 queries. This is indicated in Figure 5, which shows prominent spikes at zero repetitions and at 10 repetitions respectively for all models. Further, 58% of the time, a hallucinated package is repeated more than once in 10 iterations, which shows that a majority of hallucinations are not simply random errors, but a repeatable phenomenon that persists across multiple iterations. This is significant because a persistent hallucination is more valuable for malicious actors looking to exploit this vulnerability and makes the hallucination attack vector a more viable threat.
@@
Verbose Models vs. Conservative Models. LLMs operate with inherent randomness and uncertainty. This behavior enables novel and creative output, a desired feature for many NLP tasks but less welcome for code generation, which requires a high degree of accuracy and must adhere to rigid syntax. We define a verbose model as one that operates with higher degree of uncertainty and randomness by generating a greater number of distinct package names while a conservative model generates a lesser number of distinct packages, generally using only the most popular and well-known packages. To this end, we investigated whether verbose models correspond to higher rate of package hallucinations. Our results show a correlation between the hallucination rate and number of unique packages that were recommended during this experiment (i.e. a more verbose model was associated with a higher hallucination rate). In light of these findings, as shown in Figure 6, it is reasonable to suggest that coding models, or models performing coding tasks, should operate in a more conservative manner when generating packages to answer a coding prompt.
@@
The most successful models in our study (i.e., the lowest hallucination rates) adhered to a smaller subset of well-known packages when generating code, and these models (e.g., the GPT series) also scored the highest on the EvalPlus [45] code quality benchmarks. This suggests that improving code quality and reducing hallucinations can potentially be achieved simultaneously without a trade-off.
LLMs’ Ability to Detect Hallucinations. We then evaluated each model’s ability to identify hallucinations vs. valid packages, both from its own code generation outputs and those generated by other models. To test this, we conducted two binary classification tests: (i) each model’s ability to detect hallucinations vs. valid packages from its own generated code, and (ii) each model’s ability to detect the same from code generated by other models. Names of valid and hallucinated packages produced by each model were randomly sampled and each model was asked if the provided name was a valid Python package. The identification accuracy was calculated as the ratio of correct identifications to the total number of provided packages.
@@
Curiously, 3 of the 4 models (GPT 4 Turbo, GPT 3.5, and DeepSeek) proved to be highly adept at detecting their own hallucinations (see Figure 7) with detection accuracy above 75%. The precision and recall values for this test, which also averaged above 75%, are presented in Table 2. This phenomenon implies that each model’s specific error patterns are detectable by the same mechanisms that generate them, suggesting an inherent self-regulatory capability. The indication that these models have an implicit understanding of their own generative patterns which could be leveraged for self-improvement is an important finding for developing mitigation strategies.
@@
CodeLlama displays unique and interesting behavior during both the tests, as it has an overwhelming propensity to label most packages as valid, resulting in a lower accuracy for hallucinated packages.
RQ3 Summary: Package hallucinations are often persistently generated. Models that generate fewer packages when prompted are correlated with a reduced hallucination rate. Several models were able to accurately detect their own hallucinations.

### v3

@@ 4.2 Code Generation
Model Selection. For our experiments, we chose the models that were the highest ranked on the EvalPlus leaderboard (as of January 20, 2024) [40]. During the creation of our model list, we ignored the fine-tuned versions that were ranked below their corresponding foundational models and only selected one fine-tuned version of the same foundational model of the same parameter size [40]. EvalPlus maintains a ranking of the top performing LLMs for code correctness according to a rigorous code synthesis evaluation framework. Our goal was to include a mix of top-performing base models and a few of the best-performing fine-tuned variants. We also included the GPT series of models (GPT-3.5, GPT-4, and GPT-4 Turbo) in our experiments, which currently hold the top rankings on the leaderboard. GPT models are widely considered as SOTA in terms of code generation models at the time of writing and add value to our experiments as representative commercial models. The models were not modified or altered in any way prior to testing; they are strictly “off-the-shelf.” Table 1 provides a complete list of the models that we tested in our experiments.
@@
Language Selection. In our experiments, we focus on two of the most popular programming languages, JavaScript and Python. These languages were chosen due to their overall popularity (#1 and #2 according to the GitHub 2023 Octoverse report [12]) and their dependence on open-source repositories for package management. Other popular programming languages like Java, C, or C++ do not rely on a centralized open-source repository, as Python and JavaScript do, which is a key component of this vulnerability. The open-source package repositories for these languages, npm and PyPI, represent ecosystems of 5.1 million and 573 thousand packages, respectively [58]. Of the 16 total models tested (see Table 1), 14 were tested for both Python and JavaScript, while two fine-tuned Python-specific models, WizardCoder-Python and CodeLlama-Python, were only tested for Python.
Testing Environment. All open-source models were tested using the Hugging Face transformers package and quantized versions of the models, which reduces parameter precision to boost inference speed and lower memory use without significantly impacting performance. Specifically, the GPTQ quantization method was used, which utilizes a one-shot weight quantization method based on approximate second-order information that has a negligible effect on the accuracy of models, making it an ideal choice [10]. Additionally, quantized models better simulate the performance that a typical user can expect when running models on commercial grade hardware, making them more accessible and practical for everyday use.
For testing uniformity, we use the same parameters and quantization precision for all open-source models, which are summarized in Appendix C, along with the computing environment used. To generate code for our analysis, we query each LLM (Table 1) with prompts from the two datasets along with a system message which contains specific instructions regarding the task and output format. An overview of the process, including the system messages used during each step, is detailed in Appendix B. The experiment generates 19,200 code samples per model (16 Python tests + 14 JavaScript tests * 19,200 = 576,000 total code samples), which are further analyzed to determine which packages are required to execute the generated code.
@@ 4.3 Detection Methodology and Heuristics
To detect hallucinated packages, we first need to extract package names from the LLM output or the generated code sample, which is non-trivial. Simply parsing the code for “import” or “require” is not useful, as the arguments in those statements refer to modules and not packages. There is no way to definitively determine the required packages from a code snippet alone. A detailed explanation of this problem can be found in Appendix G. To solve this problem, we employ the following three heuristics to determine/identify package names in the generated code:
@@
Heuristic 1. As part of our first heuristic, we parse the generated Python and JavaScript code for “pip install” and “npm install” commands, respectively. These commands look for the specified package in the PyPI/npm repository, resolve its dependencies, and install everything in the current Python/JavaScript environment to ensure that future module requests will work. This is the most straightforward heuristic for detecting package names (and thus hallucinations), as it involves explicit commands from the code generation model for package download/installation. This is significant because if the referenced hallucinated package was indeed used by an adversary to execute a package confusion attack, it could immediately trigger download/install of the malicious code in the package. Note that we did not directly ask the model to provide these commands, but allowed them to occur naturally during the generation process. As such, we observed that these instances (“pip install” and “npm install”) occur for $7\%$ of the total output.
@@
Heuristic 2. For the second heuristic, each generated code sample is used as input to the same model that generated it. The model is then prompted for a list of packages that would be required to run the given code. Our intuition is to mimic an actual user/developer who is using LLMs for code generation. If the user gets an error due to an uninstalled package when attempting to execute the generated code, they could query the model for the correct package to install. We wanted to replicate this intuitive process to identify the package names required by the generated code.
@@
Heuristic 3. As the third heuristic, we reuse the original prompt used to generate the code sample as an input to the model and ask the model to output package names that would be required to accomplish this coding task. Similarly to the previous heuristic, this process of extracting package names simulates another approach users would take to obtain package names from the model that generated the code, if the required packages were not mentioned in explicit “pip install” and “npm install” commands.
@@
Once each model provides specific package names (through the three heuristics outlined above), we simply compare each package name to a master list of package names acquired from PyPI and npm, respectively (each list is as of 10 January, 2024). If a package name is not on the master list, it is considered a hallucination. We acknowledge the possibility that the master list of packages obtained from the package repositories has already been contaminated with malicious hallucinated packages. It is not possible to guarantee that the master list actually represents the ground truth of valid packages; however, the presence of hallucinated packages already in the master list would actually produce fewer hallucinations, and therefore our results represent a lower bound of hallucination rate.
@@ 5 Evaluation Results
In this section, we present the results of our experimental analysis related to RQ1 -- RQ4. After using both Python and JavaScript for RQ1, for RQs 2 through 4, we focus our analysis only on the Python programming language, a subset of the original models tested, and randomly sampled subsets of our original datasets. Given the consistent results that we were able to obtain across both languages for RQ1, we believe that this narrowed scope of discussion for RQs 2 through 4 should not compromise the generalizability of the conclusions and would allow for a deeper analysis of package hallucinations in a controlled setting. We selected GPT-4 Turbo, GPT-3.5, CodeLlama 7B, and DeepSeek 6.7B for the in-depth analysis of RQs 2-4, representing the best-performing and most popular open-source models.
@@ 5.1 Prevalence of Package Hallucinations (RQ1)
In our first experiment, our goal was to quantify the prevalence of package hallucinations across different models by generating and analyzing a large number of code samples. We conducted 30 tests (using 16 models for Python and 14 models for JavaScript, as described in Table 1) producing a combined 576,000 code samples using both the Stack Overflow and LLM-generated datasets (Section 4.1). Each code sample was evaluated for hallucinations according to the heuristics defined in Section 4.3, which include parsing the generated code and prompting the model for packages twice per code sample, for a total of 1,152,000 package prompts across all tests. To measure LLMs’ propensity to produce hallucinated packages during code generation, we use the package hallucination rate metric, which can be expressed as a simple ratio of the number of hallucinated packages to the total number of recommended packages. The total hallucination rates for each evaluated model are presented in Figure 2. More fine-grained results on hallucination rates for all models tested, covering both Python and JavaScript, are presented in Appendix E.
@@
These 30 tests generated a total of 2.23 million packages in response to our prompts, of which 440,445 (19.7%) were determined to be hallucinations, including 205,474 unique non-existent packages (i.e. packages that do not exist in PyPI or npm repositories and were distinct entries in the hallucination count, irrespective of their multiple occurrences). Our results for GPT-3.5 (5.76%) and GPT-4 (4.05%) differ significantly from previous work on package hallucinations [30], which found hallucination rates 4--6 times higher (24.2% and 22.2%, respectively) for those specific models. GPT series models were found to be 4 times less likely to generate hallucinated packages compared to open-source models, with a hallucination rate of 5.2% compared to 21.7%. GPT-4 Turbo resulted in the lowest overall hallucination rate at 3.59%, while DeepSeek 1B had the best hallucination rate among open-source models at 13.63%. Python code resulted in fewer hallucinations than JavaScript (15.8% on average compared to 21.3% for JavaScript). Despite the difference in hallucination rate between the two languages, there is a linear relationship between the results (as shown in Figure 14 in the appendix), demonstrating that the propensity of a model to hallucinate is positively correlated between programming languages. The above results provide strong evidence that package hallucinations are a pervasive issue across all code-generating LLMs.
RQ1 Summary: Package hallucinations were found to be pervasive phenomenon across all 16 models tested. Commercial models hallucinated 4$\times$ less compared to open-source models. Python code resulted in a lower hallucination rate compared to JavaScript.
@@ 5.2 Impact of Model Settings (RQ2)
Effect of Temperature Settings. The temperature setting in a LLM is used to adjust the randomness of the generated responses, where a lower temperature results in more predictable and deterministic outputs, while a higher temperature increases creativity and diversity in the responses (Section F.1). We varied this setting for each model between the minimum and maximum allowed values and observed the change in hallucination rate (the maximum temperature for the GPT series models is limited to 2, while the open-source models can be set to 5). All models exhibited a clear increase in hallucination rate as temperature value increases, with the effect becoming severe at maximum values. The OpenAI models, as shown in Figure 3, showed only a slight increase in hallucination rate between temperatures 0 and 1, which then increased sharply between 1 and 2. In particular, GPT-4 resulted in a hallucination rate (8.9%) nearly 4 times lower than GPT 3.5 (31.8%) at its maximum temperature. At the highest temperature values, open-source models start to generate more hallucinated packages than valid packages. Most LLMs operate at a default temperature in the range of 0.7 to 1, however, our results indicate that a lower temperature value can reduce package hallucinations, with the optimal value varying per model. Lower temperature also yields more deterministic responses, presenting a trade-off between risk of hallucination and creativity. Therefore, selecting the appropriate temperature value is not a straightforward decision.
@@
Effect of Decoding Strategies. Next, we adjusted several decoding parameters (top-$p$, top-$k$, and min-$p$ values) to reduce the chances of a low probability token being selected as a potential package, with the intuition that lower probability tokens correspond to higher probabilities of hallucination in this context. The following is a summary of the parameters and values we modified.
@@
We evaluated each listed value in isolation, followed by a combined evaluation of the values highlighted in bold, resulting in a total of 10 tests. Note that top-$k$ and min-$p$ were only tested for DeepSeek and CodeLlama, as these values are not modifiable through the OpenAI API. Varying the decoding values induced a slight increase (1.16% on average) in the hallucination rate for the four models across all values tested. As we will expand on in RQ3, package hallucinations are often persistently repeated across many iterations.
@@
Recency of Subject Matter. As described in Section 4.1, we separated our coding prompts into two temporal datasets to evaluate whether the hallucination rate was correlated with topics/packages that emerged after the model was trained. A lower difference between the rates of recent and all-time prompts indicates better performance in handling questions that fall outside the model’s pre-training data and therefore a more generalizable model. The models we tested were shown to be more likely to generate a package hallucination when responding to prompts that deal with more recent topics. This difference resulted in a 10% higher hallucination rate on average for older data versus more recent data.
Overall, all 16 Python models we evaluated demonstrated a higher hallucination rate when being prompted about questions or packages that were popular within the past year (Figure 4). These higher rates are at least partially due to the inherent limitations and training costs of modern LLMs. As noted in the OpenAI GPT-4 technical report [1], LLMs cannot update themselves with new information after release and have no knowledge of the world beyond their training data cutoff date. Although fine-tuning can enhance specific tasks, it generally does not improve the model’s overall knowledge of the world. The massive cost of training LLMs from scratch, evidenced by the 1,400,000 GPU hours (220 years) required to train the 12 CodeLlama models, makes continuously updating pre-training data impractically expensive [53]. This cost, along with steadily increasing model sizes and training times, poses a significant barrier to reducing package hallucinations for advanced coding prompts and packages.
@@ 5.3 Model Behaviors (RQ3)
Frequency of Repeated Hallucinations. To determine whether hallucinations are random error or repeatable phenomena, this test focuses on the persistence of hallucinations within a model. We randomly sampled 500 prompts that generated package hallucinations during our initial testing and then repeated those queries 10 times per prompt. Of those 10 queries, we recorded how many times the original hallucinated package was regenerated. Our analysis reveals an unexpected dichotomy when repeatedly querying a model with the same prompt that generated a hallucination: 43% of hallucinated packages were repeated in all 10 queries, while 39% did not repeat at all across the 10 queries. This is indicated in Figure 5, which shows prominent spikes at zero repetitions and at 10 repetitions, respectively, for all models. In addition, 58% of the time, a hallucinated package is repeated more than once in 10 iterations, which shows that a majority of hallucinations are not simply random errors, but a repeatable phenomenon that persists across multiple iterations. This is significant because a persistent hallucination is more valuable for malicious actors looking to exploit this vulnerability and makes the hallucination attack vector a more viable threat.
@@
Verbose Models versus Conservative Models. LLMs operate with inherent randomness and uncertainty. This behavior enables novel and creative output, a desired feature for many NLP tasks but less welcome for code generation, which requires a high degree of accuracy and must adhere to rigid syntax. We define a verbose model as one that operates with higher degree of uncertainty and randomness by generating a greater number of distinct package names while a conservative model generates a lesser number of distinct packages, generally using only the most popular and well-known packages. To this end, we investigated whether verbose models correspond to a higher rate of package hallucinations. Our results (Figure 6) show a correlation between hallucination rate and number of unique packages that were recommended during this experiment (i.e. a more verbose model was associated with a higher hallucination rate). In light of these findings, it is reasonable to suggest that models generating code should adopt a more conservative approach (i.e. limiting package suggestions to a smaller list of well-known packages rather than generating names with uncertainty).
@@
The models with the lowest hallucination rates in our study adhered to a smaller subset of packages when generating code, and these models (e.g., the GPT series) also scored the highest on the EvalPlus [40] code quality benchmarks. This suggests that improving code quality and reducing hallucinations can potentially be achieved simultaneously without a trade-off.
LLMs’ Ability to Detect Hallucinations. We then evaluated each model’s ability to identify hallucinations versus valid packages, both from its own code generation outputs and those generated by other models. To test this, we conducted two binary classification tests: (i) each model’s ability to detect hallucinated packages from its own generated code and (ii) each model’s ability to detect hallucinated packages from code generated by other models. The names of the valid and hallucinated packages produced by each model were randomly sampled, and each model was asked "Is [package name] a valid Python package?" Identification accuracy was calculated as the ratio of correct identifications to the total number of packages provided.
@@
Figure 7 shows that 3 of the 4 models (GPT 4 Turbo, GPT 3.5, and DeepSeek) proved to be highly adept in detecting their own hallucinations with detection accuracy above 75%. Table 2 displays the recall and precision values for this test, with similarly strong performance across the 3 proficient models. This phenomenon implies that each model’s specific error patterns are detectable by the same mechanisms that generate them, suggesting an inherent self-regulatory capability. The indication that these models have an implicit understanding of their own generative patterns that could be leveraged for self-improvement is an important finding for developing mitigation strategies.
@@
CodeLlama displays unique and interesting behavior during both tests, as it has an overwhelming propensity to label most packages as valid, resulting in a lower accuracy for hallucinated packages.
RQ3 Summary: Package hallucinations are often persistently generated. Models that generate fewer packages when prompted are correlated with a reduced hallucination rate. Several models were able to detect their own hallucinations with greater than 75% accuracy.